if four times the supplement of an angle is added to eight times the angle's complement, the sum is equivalant to three straight angles. Find the measure of the angle that is supplementry to the complement.
Answer by cleomenius(959)

4(180 - x) + 8(90 - x) = 540
720 - 4x + 720 - 8x = 540
1440 = 12x + 540
900 = 12x
x = 75 degrees. This is our angle.
The complement is 15 degrees.
The supplement of the complement is 165 degrees.
=====================================================
To check, 4(180 - x) + 8(90 - x) = 540
4(105) + 8(15) = 540
420 + 120 = 540
=====================================================
This does check.
Cleomenius.